# Jailbreak Attacks and Defenses against Multimodal Generative Models: A Survey

### v1

@@ I Introduction
The growing deployment of multimodal generative models has raised significant concerns about their security and reliability. Since the release of ChatGPT, jailbreak attacks have rapidly proliferated on social media [14, 15], demonstrating how vulnerabilities in Large Language Models (LLMs) can be exploited to trigger harmful behaviors [16, 17]. Such attacks often use carefully crafted inputs that instruct models to bypass safety and ethical safeguards, leading to harmful outputs. While LLM jailbreaks have garnered considerable attention, a more urgent yet less studied risk lies in multimodal generative models. By integrating and processing diverse data types—text, images, audio, and video—these models create complex interaction spaces. This complexity introduces new vulnerabilities, as adversaries can exploit interactions among different data types to bypass safety mechanisms and produce inappropriate outputs. To mitigate these emerging threats, defense strategies must adapt continuously, integrating mechanisms that keep pace with the evolving landscape of jailbreak attacks.
@@
Figure 2: Jailbreak attack and defense against multimodal generative models.
Existing reviews on jailbreak methods [22, 23, 24, 25] have primarily concentrated on content output within a specific modality, addressing tasks either Any-to-Text [26, 23] or Any-to-Image [24, 25]. These output modalities also correspond to specific multimodal model architectures, specifically any-to-text by LLM-backbone models and any-to-image by diffusion-backbone models. While these surveys have made valuable contributions, they lack a unified framework that spans a broad range of modalities (i.e., text, images, audio, and video) within different generative systems, as shown in Fig. 1.
@@ II-A Multimodal Generative Models
$\bullet$ Any-to-Vision Models. Diffusion models represent a major breakthrough in visual generation, surpassing traditional methods like Generative Adversarial Networks (GANs) [45]. They are extensively used in generative tasks. By operating as parameterized Markov chains, diffusion models transform random noise into complex images through sequential refinement. The reverse process is modeled as a normal distribution. A vision processor, often UNet, is trained to estimate the mean and variance.
For image generation, Ho et al.[46] introduce classifier-free guidance by training an additional text encoder as a condition interpreter, allowing for flexible conditioning with free-form text prompts. Beyond text, ControlNet[47] incorporates specific conditions like depth maps, normal maps, canny edges, poses, and sketches to enhance fidelity and controllability. Methods such as UniControlNet [48], UniControl [49], Composer [50], and T2I-Adapter [51] further improve generation by integrating diverse conditions and additional layers as condition interpreters.
Text-to-video generation models rely on three core components: condition interpreters, vision processors, and temporal handlers [52]. Condition interpreters translate the input text into visual elements, connecting the semantics of the text with the objects in the image and their dynamics in the video. Vision processors handle the visual content within each frame. Since a video is essentially a sequence of images, these models often use vision processing modules similar to those in image generation. Temporal handlers manage the progression between frames, learning the dynamics of visual content over time. This element is unique to text-to-video models, enabling the capture of motion and transitions between frames through various mechanisms.
@@
TABLE II: Notations related to jailbreak against multimodal models. For input/output modalities, I: Image, T: Text, V: Video, A: Audio.
@@
$\mathcal{X}_{mal}$ Malicious input.
$\mathcal{X}_{adv}$ Adversarial input.
$\mathcal{Y}_{mal}$ Malicious output.
$\mathcal{Y}_{adv}$ Adversarial output.
@@
$\epsilon_{\theta}\left(z_{t}|c\right)$ Predicted noise generated by Denoising Networks.
@@ II-B Jailbreak Attack
where $\mathcal{L}_{harm}$ quantifies the degree of harmfulness in the generated content $\mathcal{Y}_{mal}$, and $\mathcal{L}_{tox}$ assesses the manifest toxicity of the adversarial prompt. $\epsilon$ denotes the corresponding thresholds. The objective is to optimize the generation of highly harmful content while maintaining manifest toxicity below a certain threshold, ensuring it avoids detection and filtering.
$\bullet$ Multimodal Jailbreak. Research on jailbreak attacks against generative models has emerged from the field of LLMs [17, 53]. For instance, Zou et al. [54] proposes to append specific adversarial suffixes to malicious prompts, while AutoDAN [55] employs a sophisticated hierarchical genetic algorithm to autonomously generate subtle jailbreak prompts. As generative models evolve beyond text to include multimodal architectures such as vision-language and text-to-image models, the scope of jailbreak attacks has similarly expanded. In these settings, adversaries exploit not only textual prompts but also visual and multimodal embeddings, requiring more advanced defense mechanisms to ensure the ethical and secure use of these technologies. Continued research in this area is crucial to developing effective safeguards against adversarial manipulation while preserving the utility and accessibility of generative models across a wide range of applications.
@@ II-C Jailbreak Defense
To address these vulnerabilities, significant efforts have focused on enhancing the safety alignment of LLMs by improving data quality during the Supervised Fine-tuning (SFT) phase [56] and aligning model behavior with human safety preferences during the Reinforcement Learning from Human Feedback (RLHF) stage [57]. Although LLMs have made significant progress in safety alignment, the introduction of additional modalities in multimodal generative models adds layers of complexity and exposes new vulnerabilities. For instance, while harmful text prompts alone may be blocked, malicious content embedded as typographic images can still bypass defenses and lead to the generation of unsafe content. This underscores the importance of robust jailbreak defense mechanisms tailored to the unique structure and applications of multimodal generative models.
@@ III-A Black-box Jailbreak
In black-box scenarios, attackers lack direct access to the internal architecture, parameter configurations, or gradient details of the target model, restricting their interactions to observing and manipulating only the model’s input-output behavior.
Figure 3: Illustration of black box jailbreak attacks against multimodal generative models at the input level where attackers focus on devising sophisticated jailbreak input templates.
TABLE III: Summary of jailbreak attack methods against multimodal models. Black means black-box attack, white means white-box attack, and gray means gray-box attack. For input/output modalities, I: Image, T: Text, V: Video, A: Audio.
@@
SASP [58] [arXiv’23] Black Input Prompt Engineering I + T → T Extract the system prompts from GPT-4V for jailbreaking MLLMs.
@@
AIAH [62] [arXiv’24] Black Input Prompt Engineering A + T → T Decompose harmful words into letters and conceal them in audio input.
@@
APGP [66] [arXiv’24] Black Input Prompt Engineering T → I Use LLMs to refine prompts via score function for copyright infringement.
@@
Atlas [68] [arXiv’24] Black Input Prompt Engineering T → I Develop three LLM agents with COT reasoning for T2I jailbreak.
@@
Figstep [71] [arXiv’23] Black Input Image Engineering I + T → T Transform malicious text prompts into image form overlaid on white backgrounds.
MM-SafetyBench [72] [ECCV’24] Black Input Image Engineering I + T → T Use Stable Diffusion to generate images based on the extracted keywords.
Logic Jailbreak [73] [arXiv’24] Black Input Image Engineering I + T → T Design flowchart images to access MLLM’ logical reasoning abilities.
Visual-RolePlay [74] [arXiv’24] Black Input Role Play I + T → T Act as high-risk roles in image inputs to generate harmful content.
Voice Jailbreak [21] [arXiv’24] Black Input Role Play A → A Use fictional storytelling elements in voice prompts to jailbreak GPT-4o.
@@
Redteaming Attack [77] [ECCV’24] White Encoder Gradient-based I + T → T Mislead VLLM outputs by attacking on the vision encoder.
MMA-Diffusion [19] [CVPR’24] White Encoder Gradient-based T → I Generate adversarial prompts via gradient optimization and word regularization.
JPA [78] [arXiv’24] White Encoder Gradient-based T → I Optimize prefix prompts in latent space to align malicious concepts.
@@
RT-Attack [80] [arXiv’24] Gray Encoder Search-based T → I Two-stage random search to optimize prompt-wise and image-wise similarity.
Image Hijacks [81] [arXiv’24] White Generator Gradient-based I + T → T Optimize image hijacks using behavior matching for multi-type attacks.
@@
ImgJP [83] [arXiv’24] White Generator Gradient-based I + T → T Generate adversarial images based on maximum likelihood algorithm.
ImgJS [84] [arXiv’24] White Generator Gradient-based I + T → T Convert visual adversarial vectors to text space, merged with harmful queries.
@@
UMK [88] [arXiv’24] White Generator Gradient-based I + T → T Optimize adversarial prefixed across image-text modalities simultaneously.
@@ A.1 Input-level Attack
As shown in Fig. 3, attackers are compelled to develop more sophisticated input templates across prompt engineering, image engineering, and role-ploy techniques. These techniques can bypass the model’s safeguards, making the models more susceptible to executing prohibited instructions.
@@
Prompt Engineering. Prompt engineering involves strategically modifying specific words or rewriting complete prompts. We broadly divide this approach into three types based on prompt sources: manual construction, generation using off-the-shelf LLM models, and generation using trained proxy models.
@@
$\bullet$ Image + Text → Text (IT2T) Models. SASP [58] employs carefully crafted dialogues to extract confidential system prompts, facilitating the discovery of potential jailbreak prompts. Furthermore, these attack prompts are incorporated with manual modifications to enhance the attack success rate. Compared with manual construction, AutoJailbreak [59] leverages LLMs as red-team tools to automatically refine attack prompts. This framework integrates weak-to-strong prompt optimization strategies, a suffix-based attack enhancement technique, and an efficient search mechanism. Instead of relying on off-the-shelf LLMs, other methods involve optimizing proxy models for prompt refinement. Arondight [60] automates multimodal jailbreak attacks by generating textual prompts using a trained LLM, supplemented with toxic images crafted by the proprietary GPT-4V. The red team LLM, guided by a reinforcement learning agent, is optimized by incorporating entropy bonuses and novelty reward metrics to enhance prompt diversity. Similarly, AdvWeb [61] proposes the first black-box framework specifically designed to mislead MLLM-powered web agents. This framework optimizes a generative model to produce adversarial prompts, which are injected into web pages to execute targeted malicious actions.
$\bullet$ Audio + Text → Text (AT2T) Models. AIAH [62] introduces a speech-specific jailbreaking technique, that breaks down harmful words into individual letters to obscure their presence in the audio input. The model is then instructed to concatenate these letters back into complete words, thereby reconstructing the original harmful question within the jailbreak prompt.
$\bullet$ Text → Image (T2I) Models. Earlier work [63] proposes a “prompt dilution” strategy to degrade filter performance by adding an unrelated extra to prompts. Through reverse engineering, it is shown that the filter specifically targets sexual content while ignoring other problematic themes like violence.
@@
Advances in off-the-shelf LLMs have also reduced the barrier to producing diverse jailbreak prompts on T2I models. By utilizing the concept of “substitution”, both PGJ [64] and SurrogatePrompt [65] employ LLMs, such as ChatGPT, to replace unsafe content with benign phrases. These approaches typically introduce perceptual confusion by preserving visual similarity while altering the textual semantics (e.g., replacing “blood” with “watermelon juice”). Unlike replacement-based methods, the automated Prompt Generation Pipeline (APGP) [66] utilizes LLMs to refine complete jailbreaking prompts by optimizing a self-generated score and imposing penalties for certain keywords, thereby effectively assessing copyright infringement risks in T2I systems. Furthermore, agent-based methods can unleash the potential of LLMs in prompt engineering. Divide-and-Conquer Attack (DACA) [67] leverages GPT models as text transformation agents to create adversarial prompts. They design attack helper prompts that guide LLMs to break down an unethical drawing intent into multiple benign descriptions of individual image elements. Building on a collaborative multi-agent framework, Atlas [68] employs two specialized agents: the mutation agent and the selection agent. Each agent is equipped with four modules: a MLLM/LLM brain, planning, memory, and tool usage. These two agents collaborate in an iterative process to determine jailbreak prompt-triggered responses and generate new candidate jailbreak prompts.
@@
In a similar vein, training proxy models can support the development of specialized red-team tools for jailbreak-prompt generation in T2I models. The Unified Prompt Attack Model (UPAM)[69] aims to optimize LoRA adapters within LLMs to generate natural adversarial prompts by utilizing a two-stage learning framework. In the first stage, sphere-probing learning optimizes prompts to bypass defenses, while the second stage employs semantic-enhancing learning to refine the generated images to align with harmful targets. Additionally, BPSA[70] optimizes a text retriever to identify sensitive words associated with the input vector and leverages an LLM API to facilitate the generation of stealthy attack prompts.
@@
Image Engineering. Image engineering methods primarily embed toxic content into carefully crafted images by text typography and diffusion-based generation. Due to the insufficient security training in the visual modules of current generative models [71], it is relatively more vulnerable to jailbreak attacks.
@@
$\bullet$ Image + Text → Text (IT2T) Models. Figstep [71] proposes to transform malicious text prompts into image-based typographic forms. This approach converts toxic questions into textual statements and then overlays these statements onto a white background image, along with indices as cues for response formatting. These images are subsequently paired with benign text prompts to execute the jailbreak attack. Some other methods have explored the potential of using diffusion models to generate targeted jailbreak images. MM-SafetyBench [72] employs a dual-approach strategy to generate images from malicious text prompts. The first approach utilizes stable diffusion [10] to generate images based on the extracted keywords from the malicious prompts. The second approach converts the keywords or entities into visual typographic forms. Additionally, the Logical jailbreak method [73] designs flowchart images with semantic meaning to evaluate the logical reasoning and visual imagination capabilities of MLLMs.
Role Play. Role-play methods often construct targeted scenarios that induce the model to take on a role aligned with harmful behaviors, leading to successful jailbreaking attempts. This method leverages the model’s capacity for complex interactions and role simulation, often through visually-driven or story-driven roles.
$\bullet$ Image + Text → Text (IT2T) Models. Visual-RolePlay [74] introduces a role-playing framework that generates images depicting high-risk roles, with descriptive text positioned at the top and malicious prompts at the bottom. After generating these role-specific images, the model is instructed to play high-risk roles in image inputs to produce harmful content.
$\bullet$ Audio → Audio (A2A) Models. VoiceJailbreak [21] introduces an innovative voice-based jailbreak attack on GPT-4o by exploiting fictional storytelling techniques. This method constructs prompts around key elements of fictional writing – such as setting, character, and plot. By engaging the model as a reader of fictional narratives, the adversary rephrases forbidden questions as assertive statements within fictional contexts, achieving a higher attack success rate through multi-step voice inputs.

### v2

@@ I Introduction
The growing deployment of multimodal generative models has raised significant concerns about their security and reliability. Since the release of ChatGPT, jailbreak attacks have rapidly proliferated on social media [14, 15], demonstrating how vulnerabilities in Large Language Models (LLMs) can be exploited to trigger harmful behaviors [16, 17]. Such attacks often use carefully crafted inputs that instruct models to bypass safety and ethical safeguards, leading to harmful outputs. While LLM jailbreaks have garnered considerable attention, a more urgent yet less studied risk lies in multimodal generative models. By integrating and processing diverse data types (i.e., text, images, audio, and video), these models create complex interaction spaces. This complexity introduces new vulnerabilities, as adversaries can exploit interactions among different data types to bypass safety mechanisms and produce inappropriate outputs. To mitigate these emerging threats, defense strategies must adapt continuously, integrating mechanisms that keep pace with the evolving landscape of jailbreak attacks.
@@
Figure 2: Jailbreak attacks and defenses against multimodal generative models. Given the generalized lifecycle of multimodal jailbreak, we systematically explore attacks and defense strategies across four levels: input, encoder, generator, and output.
Existing reviews on jailbreak methods [22, 23, 24, 25] have primarily concentrated on content output within a specific modality, addressing tasks either Any-to-Text [22, 26, 23] or Any-to-Image [24, 25]. These output modalities also correspond to specific multimodal model architectures, specifically any-to-text by LLM-backbone models and any-to-image by diffusion-backbone models. While these surveys have made valuable contributions, they lack a unified framework that spans a broad range of modalities (i.e., text, images, audio, and video) within different generative systems, as shown in Fig. 1.
@@ II-A Multimodal Generative Models
$\bullet$ Any-to-Vision Models. Diffusion models represent a major breakthrough in visual generation, surpassing traditional methods like Generative Adversarial Networks (GANs) [45]. Specifically, diffusion models treat image generation as a parameterized Markov chain. They generate new images by employing a backward process that iteratively denoises Gaussian noise $z_{T}$. A parameterized Gaussian transition network is introduced to model this backward process. In practice, a vision processor, typically a UNet, is trained to estimate the mean and variance of the Gaussian transition network. For controllable image generation, an additional text encoder is introduced as a condition interpreter, allowing for flexible conditioning with free-form text prompts [46, 33, 34].
Text-to-video generation models rely on three core components: condition interpreters, vision processors, and temporal handlers [35, 36]. Condition interpreters translate the input text into visual elements, connecting the semantics of the text with the objects in the image and their dynamics in the video. Vision processors handle the visual content within each frame. Since a video is essentially a sequence of images, these models often use vision processing modules similar to those in image generation. Temporal handlers manage the progression between frames, learning the dynamics of visual content over time. This element is unique to video generation models, enabling the capture of motion and transitions between frames through various mechanisms.
@@
TABLE II: Notations related to jailbreak attacks and defenses against multimodal generative models.
@@
$x_{mal}$ Malicious input.
$x_{adv}$ Adversarial input.
$c_{mal}$ Malicious concept.
$y_{mal}$ Malicious output.
$y_{adv}$ Adversarial output.
@@
$\epsilon_{\theta}\left(z_{t},e_{x},t\right)$ Predicted noise generated by Denoising Networks.
@@ II-B Jailbreak Attack
where $\mathcal{S}_{harm}$ quantifies the degree of harmfulness in the generated content, and $\mathcal{S}_{tox}$ assesses the manifest toxicity of the adversarial prompt. $\epsilon$ denotes the corresponding thresholds. The objective is to optimize the generation of highly harmful content while maintaining manifest toxicity below a certain threshold, ensuring it avoids detection and filtering.
$\bullet$ Multimodal Jailbreak. Research on jailbreak attacks has emerged from the field of LLMs [17, 47, 48, 49, 50]. For instance, Zou et al. [48] propose to append specific adversarial suffixes to malicious prompts, while AutoDAN [51] employs a sophisticated hierarchical genetic algorithm to autonomously generate subtle jailbreak prompts. As generative models evolve beyond text to include multimodal architectures such as vision-language and text-to-image models, the scope of jailbreak attacks has similarly expanded. In these multimodal settings, attackers can exploit not only textual prompts [52, 53, 54] but also vulnerabilities in visual inputs [18, 55] and multimodal embeddings [56], targeting the complex interactions between text, images, and other modalities. This expanded attack surface presents significant challenges to ensuring the safety and robustness of multimodal generative models.
@@ II-C Jailbreak Defense
To mitigate these vulnerabilities, current efforts focus on enhancing the safety alignment of LLMs by improving data quality during the Supervised Fine-tuning (SFT) phase [57] and aligning model behavior with human safety preferences during the RLHF stage [58]. Although LLMs have made significant progress in safety alignment, the introduction of additional modalities in multimodal generative models adds layers of complexity and exposes new vulnerabilities. For instance, incorporating additional modality encoders introduces risks associated with the encoding of unsafe embeddings [59, 60, 61]. These risks can lead to the propagation of harmful content across modalities, potentially compromising the reliability of the overall generative system. This requires more advanced defense mechanisms to ensure the ethical and secure use of generative models while preserving the utility of these technologies across a wide range of applications.
@@ III-A Black-box Jailbreak
In black-box scenarios, attackers lack access to the internal architecture, parameter configurations, or gradient details of the target model, restricting their interactions to observing and manipulating only the model’s input-output behavior.
Figure 3: Illustration of black box jailbreak attacks against multimodal generative models at the input level where attackers focus on devising sophisticated jailbreak input patterns.
TABLE III: Summary of jailbreak attack methods against multimodal models. For model accessibility, black: black-box attack, white: white-box attack, and gray: gray-box attack. For input/output modalities, I: Image, T: Text, V: Video, A: Audio.
@@
ColJailBreak [68] [NeurIPS’24] Black Input Prompt Engineering T → I Use LLMs for unsafe word substitution and editing generations to embed harm.
@@
Figstep [71] [arXiv’23] Black Input Transfer Injection I + T → T Transform malicious text prompts into image form overlaid on white backgrounds.
MM-SafetyBench [55] [ECCV’24] Black Input Transfer Injection I + T → T Use Stable Diffusion to generate images based on the extracted keywords.
Logic Jailbreak [72] [arXiv’24] Black Input Transfer Injection I + T → T Design flowchart images to access MLLM’ logical reasoning abilities.
Visual-RolePlay [73] [arXiv’24] Black Input Transfer Injection I + T → T Act as high-risk roles in image inputs to generate harmful content.
AIAH [74] [arXiv’24] Black Input Transfer Injection A + T → T Decompose harmful words into letters and conceal them in audio input.
Zer0-Jack [75] [arXiv’24] Black Output Estimation-based I + T → T Estimate zero-order gradients with output logits to optimize part of the image.
@@
RT-Attack [78] [arXiv’24] Black Output Search-based T → I Two-stage random search to optimize prompt-wise and image-wise similarity.
SASP [79] [arXiv’23] Black Output Multi-turn Dialogue I + T → T Leverage stolen system prompts from GPT-4V for self-adversarial attacks.
SSA [80] [arXiv’24] Black Output Multi-turn Dialogue I + T → T Use agents and tools for response generation and harmful snowballing.
IDEATOR [81] [arXiv’24] Black Output Multi-turn Dialogue I + T → T Use VLM agents to refine attack strategy based on previous responses.
APGP [82] [arXiv’24] Black Output Multi-turn Dialogue T → I Use LLMs to iteratively revise prompts based on the score function.
CoJ [83] [arXiv’24] Black Output Multi-turn Dialogue T → I Decompose malicious queries into harmless sub-queries to iteratively edit images.
ICER [84] [arXiv’24] Black Output Multi-turn Dialogue T → I Use LLMs to learn from successful in-Context red-teaming experiences.
Atlas [85] [arXiv’24] Black Output Multi-turn Dialogue T → I Develop LLM-based multi-agents combined with ICL and CoT reasoning.
Voice Jailbreak [21] [arXiv’24] Black Output Multi-turn Dialogue A → A Use fictional storytelling elements in voice prompts to jailbreak GPT-4o.
@@
Redteaming Attack [86] [ECCV’24] Gray Encoder Gradient-based I + T → T Mislead VLLM outputs by attacking on the vision encoder.
MMA-Diffusion [19] [CVPR’24] Gray Encoder Gradient-based T → I Generate adversarial prompts via gradient optimization and word regularization.
JPA [54] [arXiv’24] Gray Encoder Gradient-based T → I Optimize prefix prompts in latent space to align malicious concepts.
@@
Image Hijacks [87] [ICML’24] White Generator Gradient-based I + T → T Optimize image hijacks using behavior matching for multi-type attacks.
@@
ImgJS [89, 90] [arXiv’24] White Generator Gradient-based I + T → T Convert visual adversarial vectors to text space, merged with harmful queries.
@@
UMK [94] [ACM MM’24] White Generator Gradient-based I + T → T Optimize adversarial prefixed across image-text modalities simultaneously.
@@
TSCO [56] [arXiv’24] White Generator Gradient-based I + T → I + T Approximates image tokenization with a continuous function.
@@ A.1 Input-level Attack
As shown in Fig. 3, attackers are compelled to develop more sophisticated input patterns across prompt engineering and transfer injection techniques. These techniques can bypass the model’s safeguards, making the models more susceptible to executing prohibited instructions.
@@
Prompt Engineering. Prompt engineering involves strategically modifying specific words or rewriting complete prompts. We broadly divide this approach into two types based on prompt generation sources: using off-the-shelf LLM models and using trained proxy models.
@@
$\bullet$ Image + Text → Text (IT2T) Models. AutoJailbreak [62] leverages LLMs as red-team tools to automatically refine attack prompts. This framework integrates weak-to-strong prompt optimization strategies, a suffix-based attack enhancement technique, and an efficient search mechanism. Instead of relying on off-the-shelf LLMs, other methods [63, 64] optimize proxy models for prompt refinement. Arondight [63] automates multimodal jailbreak attacks by generating textual prompts using a trained LLM, supplemented with toxic images crafted by the proprietary GPT-4V. The red team LLM, guided by a reinforcement learning agent, is optimized by incorporating entropy bonuses and novelty reward metrics to enhance prompt diversity. Similarly, AdvWeb [64] proposes the first black-box framework specifically designed to mislead MLLM-powered web agents. This framework optimizes a generative model to produce adversarial prompts, which are injected into web pages to execute targeted malicious actions.
@@
$\bullet$ Text → Image (T2I) Models. Earlier work [65] proposes a “prompt dilution” strategy to degrade filter performance by adding an unrelated extra to prompts. Recently, advances in off-the-shelf LLMs have reduced the barrier to producing diverse jailbreak prompts on T2I models. A group of works [66, 67, 68] introduces the concept of “substitution” by instructing LLMs to replace unsafe content with benign phrases. Specifically, PGJ [66] introduces perceptual confusion by preserving visual similarity while altering the textual semantics (e.g., replacing “blood” with “watermelon juice”). ColJailBreak [68] targets normal images generated by substituting unsafe words, and injects malicious elements into them through editing. Unlike substitution-based methods, agent-based methods can better unleash the potential of LLMs. DACA [69] leverages LLMs as text transformation agents to create adversarial prompts. They design attack helper prompts that guide LLMs to break down an unethical drawing intent into multiple benign descriptions of individual image elements.
@@
In a similar vein, training proxy models can support the development of specialized red-team tools for jailbreak-prompt generation. UPAM [70] aims to optimize LoRA adapters within LLMs to generate natural adversarial prompts by utilizing a two-stage learning framework. In the first stage, sphere-probing learning optimizes prompts to bypass defenses, while the second stage employs semantic-enhancing learning to refine the generated images to align with harmful targets. Additionally, BPSA [52] optimizes a text retriever to identify sensitive words associated with the input vector and leverages LLMs to facilitate the generation of stealthy attack prompts.
@@
Transfer Injection. Transfer Injection methods primarily transfer toxic prompts into other modalities, such as images by text typography or diffusion-based generation. Due to the insufficient security training in the visual modules of current models [71], it is relatively more vulnerable to jailbreak attacks.
@@
$\bullet$ Image + Text → Text (IT2T) Models. Some works [71, 55, 92, 73] propose to create typography-based visual prompt images by overlaying malicious textual statements onto white background images. These images are subsequently paired with benign text prompts to execute the jailbreak attack. Apart from typography-based methods, some other methods [55, 72, 73] have explored the potential of using diffusion models to generate targeted jailbreak images. For instance, MM-SafetyBench [55] generates target images based on the extracted keywords from the malicious prompts. Moreover, the Logical jailbreak method [72] designs flowchart images corresponding to harmful behaviors to evaluate the logical reasoning and visual imagination capabilities of MLLMs. To leverage the model’s capacity for role simulation, Visual-RolePlay [73] generates images depicting high-risk roles, with descriptive text positioned at the top and malicious prompts at the bottom. After generating these role-specific images, the model is instructed to play high-risk roles in image inputs to produce harmful content.
$\bullet$ Audio + Text → Text (AT2T) Models. AIAH [74] introduces a speech-specific technique, that breaks down harmful words into individual letters to obscure their presence in the audio input. The model is then instructed to concatenate these letters back into complete words, thereby reconstructing the original harmful question within the jailbreak prompt.
A.2 Output-level Attack
In Fig. 4, attackers focus on iteratively querying the model’s responses to refine inputs based on optimization-based and multi-turn dialogue methods. Optimization-based methods typically rely on specific adversarial goals, which are addressed by estimation-based and search-based attack techniques.